# Jarník-Besicovitch type theorems for semisimple algebraic groups

### v1

@@ 6. An upper bound for the Hausdorff dimension of $S_{\gamma}^{c}$
where $a=\operatorname{diag}(e^{t_{1}},e^{t_{2}},\cdots,e^{t_{n}})$. Let $\{e_{1},e_{2},...,e_{n}\}$ be the standard basis in $V$ where
@@
Suppose that $x\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ is a rational element in $\mathbf{G}(\mathbb{R})$. Let
$x=\left(\begin{array}[]{cccc}1&0&\cdots&\frac{p_{1}}{q}\\ 0&1&\cdots&\frac{p_{2}}{q}\\ \vdots&\vdots&\cdots&\vdots\\ 0&0&\cdots&1\end{array}\right)$
where $p_{i}\in\mathbb{Z},q\in\mathbb{N}$ and $(p_{1},p_{2},\dots,p_{n-1},q)=1$. Then $x\cdot\mathbb{Z}^{n}\cap V_{\beta_{0}}\neq\{0\}$ implies that there exists a primitive integral point $(k_{1},k_{2},\dots,k_{n-1},l)^{T}\in\mathbb{Z}^{n}\setminus\{0\}$ such that
$\left(\begin{array}[]{cccc}1&0&\cdots&\frac{p_{1}}{q}\\ 0&1&\cdots&\frac{p_{2}}{q}\\ \vdots&\vdots&\cdots&\vdots\\ 0&0&\cdots&1\end{array}\right)\cdot\left(\begin{array}[]{c}k_{1}\\ k_{2}\\ \vdots\\ l\end{array}\right)=\left(\begin{array}[]{c}0\\ 0\\ \vdots\\ \end{array}\right).$
Then we have
$k_{i}+\frac{p_{i}}{q}\cdot l=0,\quad 1\leq i\leq n-1.$
Hence $|l|=q$ and the denominator $d(x)$ of $x$ is equal to $q$. One can then deduce from the formula $d(x)=q$ that for any bounded open subset $\tilde{U}$ in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$ and any sufficiently large $l>0$, the number of rational elements in $\tilde{U}$ whose denominators are less than $l$ is bounded by $\mu_{R_{u}(\overline{\mathbf{P}}_{\beta_{0}})}(\tilde{U})\cdot l^{n}.$ Therefore, we may choose $A_{e}=n$ in Equation ($\ast$6) in §6. One can also compute that $a_{e}=n-1$.
Now consider the upper bound of the Hausdorff dimension of $S_{\tau}^{c}$. In the computation, we actually use the Bruhat decomposition as
$\mathbf{G}(\mathbb{R})=\bigcup_{i\in\bar{I}}\mathbf{P}_{\beta_{0}}(\mathbb{R})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{R}).$
In our case, we may choose the index set $\bar{I}$ as small as possible (as the double cosets in the original decomposition with respect to Weyl group ${}_{\mathbb{Q}}W$ overlap). Note that ${}_{\mathbb{Q}}W$ is isomorphic to the symmetric group of permutations of $\{1,2,\dots,n\}$ and many elements of ${}_{\mathbb{Q}}W$ are inside $\mathbf{P}_{\beta_{0}}(\mathbb{R})$. From this observation, one can deduce that if $w_{i}=(i,n)$ $(1\leq i\leq n)$ is the permutation of $i$ and $n$, then
$\mathbf{G}(\mathbb{R})=\bigcup_{i=1}^{n}\mathbf{P}_{\beta_{0}}(\mathbb{R})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{R}).$
Then by definition, one may compute that $A_{w_{i}}\leq A_{e}$, $a_{w_{i}}\leq a_{e}$ for all $1\leq i\leq n$, and
$\mathbf{F}_{w_{i}}=\left(\begin{array}[]{cccccc}1&0&\cdots&*&\cdots&0\\ 0&1&\cdots&*&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&1&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0&\cdots&1\end{array}\right)$
where the only non-zero entries are in the $i$-th column.
Now we compute the upper bound of $\dim_{H}S_{\tau}^{c}$. Fix $w_{i}$ $(1\leq i\leq n)$. Note that the roots in $\mathbf{F}_{w_{i}}$ are $\alpha_{1,i},\alpha_{2,i},\dots,\alpha_{i-1,i}$. According to Theorem 1.4 and the discussion in §6, we may assume that $\beta_{0}(w_{i}a_{-1}w_{i}^{-1})>0$, i.e. $b_{i}<0$; otherwise there is nothing to compute. Then $\nu_{0}=\alpha_{1,n}$ and by the fact that $b_{1}\geq b_{2}\geq\cdots\geq b_{n}$ and $b_{1}+b_{2}+\cdots+b_{n}=0$ we have
$\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w_{i}})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(w_{i}a_{-1}w_{i}^{-1})-\tau)}{\nu_{0}(a_{1})}\cdot\max\{A_{w_{i}},a_{w_{i}}\}\cdot\dim V_{\beta_{0}}$
$=$ $\dim\mathbf{G}-\frac{(b_{1}-b_{i})+(b_{2}-b_{i})+\cdots+(b_{i-1}-b_{i})}{b_{1}-b_{n}}+\frac{-b_{i}-\tau}{b_{1}-b_{n}}\cdot\max\{A_{w_{i}},a_{w_{i}}\}$
$\leq$ $\dim\mathbf{G}-\frac{(b_{1}-b_{i})+(b_{2}-b_{i})+\cdots+(b_{i-1}-b_{i})}{b_{1}-b_{n}}+n\cdot\frac{-b_{i}-\tau}{b_{1}-b_{n}}$
$=$ $\dim\mathbf{G}-\frac{(b_{1}-b_{i})+(b_{2}-b_{i})+\cdots+(b_{i-1}-b_{i})+nb_{i}}{b_{1}-b_{n}}-\frac{n\tau}{b_{1}-b_{n}}$
$=$ $\dim\mathbf{G}-\frac{-b_{i+1}\cdots-b_{n}+(n-i)b_{i}}{b_{1}-b_{n}}-\frac{n\tau}{b_{1}-b_{n}}$
$=$ $\dim\mathbf{G}-\frac{(b_{i}-b_{i+1})+\cdots(b_{i}-b_{n})}{b_{1}-b_{n}}-\frac{n\tau}{b_{1}-b_{n}}$
$\leq$ $\dim\mathbf{G}-\frac{n\tau}{b_{1}-b_{n}}.$
Therefore, by Theorem 1.4, we have
$\dim_{H}S_{\tau}^{c}\leq\dim\mathbf{G}-\frac{n\cdot\tau}{\nu_{0}(a_{1})}.$
On the other hand, for the lower bound of $\dim_{H}S_{\tau}^{c}$, by Theorem 1.3, we have
$\dim_{H}S_{\tau}^{c}\geq$ $\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\alpha(a_{1})\cdot\frac{\tau}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}$
$=$ $\dim\mathbf{G}-((b_{1}-b_{n})+(b_{2}-b_{n})+\cdots+(b_{n-1}-b_{n}))\cdot\frac{\tau}{(-b_{n})(b_{1}-b_{n})}$
$=$ $\dim\mathbf{G}-\frac{n\tau}{b_{1}-b_{n}}=\dim\mathbf{G}-\frac{n\cdot\tau}{\nu_{0}(a_{1})}.$
We conclude that
$\dim_{H}S_{\tau}^{c}=\dim\mathbf{G}-\frac{n\cdot\tau}{\nu_{0}(a_{1})}.$
This completes the proof of Theorem 1.6. ∎
Proof of Theorem 1.7.
Let $\mathbf{T}$ be the full diagonal group in $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$ and we may write
where $\mathbf{T}(\mathbb{R})^{0}$ is the connected component of identity in $\mathbf{T}(\mathbb{R})$. Let $\mathbf{P}_{0}$ be the lower triangular subgroup in $\mathbf{G}$ and $R_{u}(\overline{\mathbf{P}}_{0})$ is the upper triangular unipotent subgroup in $\mathbf{G}$. Let $V_{\beta_{0}}$ be the highest weight space in $\mathfrak{sl}_{n}$
@@
Note that here $\nu_{0}=-\beta_{0}$ is the highest root in $\mathfrak{sl}_{n}$ and $\beta_{0}(a_{t})=(b_{n}-b_{1})t<0$. The stabilizer $\mathbf{P}_{\beta_{0}}$ of $V_{\beta_{0}}$ and $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ are
@@
Let $x\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ be a rational element in $\mathbf{G}(\mathbb{R})$. We write
$x=\left(\begin{array}[]{ccccc}1&\frac{a_{1}}{b}&\cdots&\frac{a_{n-2}}{b}&0\\ 0&1&\cdots&0&0\\ \vdots&\vdots&\cdots&\vdots&\vdots\\ 0&0&\cdots&1&0\\ 0&0&\cdots&0&1\end{array}\right)\cdot\left(\begin{array}[]{ccccc}1&0&\cdots&0&\frac{p_{1}}{q}\\ 0&1&\cdots&0&\frac{p_{2}}{q}\\ \vdots&\vdots&\cdots&\vdots&\vdots\\ 0&0&\cdots&1&\frac{p_{n-1}}{q}\\ 0&0&\cdots&0&1\end{array}\right)$
where $(a_{1},a_{2},\dots,a_{n-2},b)=1$ and $(p_{1},p_{2},\dots,p_{n-1},q)=1$. Then by definition,
$x\cdot\mathfrak{sl}_{n}(\mathbb{Z})\cdot x^{-1}\cap V_{\beta_{0}}\neq\{0\}.$
It implies that there exists $l\in\mathbb{R}$ such that
$x^{-1}\cdot\left(\begin{array}[]{cccc}0&0&\cdots&0\\ \vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0\\ l&0&\cdots&0\end{array}\right)\cdot x\in\mathfrak{sl}_{n}(\mathbb{Z}).$
One can compute that
$\left(\begin{array}[]{ccccc}-\frac{p_{1}}{q}\cdot l&-\frac{p_{1}}{q}\cdot l\cdot\frac{a_{1}}{b}&\cdots&-\frac{p_{1}}{q}\cdot l\cdot\frac{a_{n-2}}{b}&-\frac{p_{1}}{q}\cdot l\cdot\frac{bp_{1}+a_{1}p_{2}+\cdots+a_{n-2}p_{n-1}}{bq}\\ \vdots&\vdots&\cdots&\vdots&\vdots\\ -\frac{p_{n-1}}{q}\cdot l&-\frac{p_{n-1}}{q}\cdot l\cdot\frac{a_{1}}{b}&\cdots&-\frac{p_{n-1}}{q}\cdot l\cdot\frac{a_{n-2}}{b}&-\frac{p_{n-1}}{q}\cdot l\cdot\frac{bp_{1}+a_{1}p_{2}+\cdots+a_{n-2}p_{n-1}}{bq}\\ l&l\cdot\frac{a_{1}}{b}&\cdots&l\cdot\frac{a_{n-2}}{b}&l\cdot\frac{bp_{1}+a_{1}p_{2}+\cdots+a_{n-2}p_{n-1}}{bq}\end{array}\right)\in\mathfrak{sl}_{n}(\mathbb{Z})$
and consequently
$\gcd(p_{1},...,p_{n-1})\cdot\frac{l}{q},\;\gcd(p_{1},...,p_{n-1})\cdot\frac{l}{q}\cdot\frac{a_{i}}{b}(1\leq i\leq n-2),\;\gcd(p_{1},...,p_{n-1})\cdot\frac{l}{q}\cdot\frac{d}{bq}\in\mathbb{Z}$
$l,\;l\cdot\frac{a_{1}}{b},\dots,l\cdot\frac{a_{n-2}}{b},\;l\cdot\frac{d}{bq}\in\mathbb{Z}$
where $d=bp_{1}+a_{1}p_{2}+\cdots+a_{n-2}p_{n-1}$. Then one can deduce from the formulas above that
$l=s\cdot\frac{bq^{2}}{\gcd(d,q)}\;(s\in\mathbb{Z}).$
Therefore, the denominator $d(x)$ of $x$ is equal to
$d(x)=\frac{bq^{2}}{\gcd(d,q).}$
It follows from the same argument as in [17, Lemma 8.6] that
|$\left|\{x\in\tilde{U}:d(x)\leq l\}\right|\ll_{\epsilon,\tilde{U}}l^{n-1+\epsilon}$
for any bounded open subset $\tilde{U}$ in $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{R})$, and hence $A_{e}=n-1$.
Now we compute the upper bound of $\dim_{H}S_{\tau}^{c}$ using Theorem 1.4. We choose elements $\{w_{i}\}_{i\in\bar{I}}$ in the Weyl group ${}_{\mathbb{Q}}W$ such that the Bruhat decomposition stated in Theorem 1.4 holds
$\mathbf{G}(\mathbb{Q})=\bigcup_{i\in\bar{I}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q}).$
Note that many Weyl elements in ${}_{\mathbb{Q}}W$ are inside $\mathbf{P}_{\beta_{0}}$. Here we choose $\{w_{i}\}_{i\in\bar{I}}$ as follows. We identify ${}_{\mathbb{Q}}W$ with the symmetric group of permutations on $\{1,2,...,n\}$. Then there are three types of elements in $\{w_{i}\}_{i\in\bar{I}}$:
Type I:
$(1,2,\dots,n-1,n)$, $(n,2,\dots,n-1,1)$
Type II:
$(1,2\dots,j-1,n,j+1,\dots,n-1,j)$, $(j,2,\dots,j-1,n,j+1,\dots,n-1,1)$, $(j,2,\dots,j-1,1,j+1,\dots,n-1,n)$, $(n,2,\dots,j-1,1,j+1,\dots,n-1,j)$ ($12\leq j\leq n-$ 1)
Type III:
$(i,2,\dots i-1,1,i+1,\dots,j-1,n,j+1,\dots,n-1,j)$, $(j,2,\dots i-1,1,i+1,\dots,j-1,n,j+1,\dots n-1,i)$ ($12\leq i<j\leq n-$ 1).
The index set $\bar{I}$ has cardinality $n(n-1)$, and one may check that $\{w_{i}\}_{i\in\bar{I}}$ meets our requirement. Note that $A_{w_{i}}\leq A_{e}$ for any $i\in\bar{I}$.
For a Weyl element $w$ of type I in $\{w_{i}\}_{i\in\bar{I}}$, if $w=(1,2,\dots,n)$ is the identity element, then the number $a_{w}$ in Corollary 6.8 comes from the morphism $\Psi_{w}$ defined by
$\left(\begin{array}[]{ccccc}1&x_{1}&\cdots&x_{n-2}&z\\ 0&1&\cdots&0&y_{1}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&1&y_{n-2}\\ 0&0&\cdots&0&1\end{array}\right)\cdot\left(\begin{array}[]{ccccc}0&0&\cdots&0&0\\ 0&0&\cdots&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0&0\\ 1&0&\cdots&0&0\end{array}\right)\cdot\left(\begin{array}[]{ccccc}1&x_{1}&\cdots&x_{n-2}&z\\ 0&1&\cdots&0&y_{1}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&1&y_{n-2}\\ 0&0&\cdots&0&1\end{array}\right)^{-1}$
$=$ $\left(\begin{array}[]{ccccc}z&-x_{1}\cdot z&\cdots&-x_{n-2}\cdot z&(x_{1}y_{1}+\cdots+x_{n-2}y_{n-2}-z)z\\ y_{1}&-x_{1}\cdot y_{1}&\cdots&-x_{n-2}\cdot y_{1}&(x_{1}y_{1}+\cdots+x_{n-2}y_{n-2}-z)y_{1}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ y_{n-2}&-x_{1}\cdot y_{n-2}&\cdots&-x_{n-2}\cdot y_{n-2}&(x_{1}y_{1}+\cdots+x_{n-2}y_{n-2}-z)y_{n-2}\\ 1&-x_{1}&\cdots&-x_{n-2}&x_{1}y_{1}+\cdots+x_{n-2}y_{n-2}-z\end{array}\right)$
where $x_{1},x_{2},\dots,x_{n-2},y_{1},y_{2},\dots,y_{n-2},z$ are the variables of $\Psi_{w}$. One can estimate $\operatorname{Vol}(E_{w}(R))$ and compute that $a_{w}\leq n-1$. Then according to Theorem 1.4, for $w$, we have

### v2

@@ 6. An upper bound for the Hausdorff dimension of $S_{\rho}(\psi)^{c}$
Let $\kappa<0$ such that
$\beta_{0}=\kappa\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}{\alpha}.$
By Lemma 7.2, one can compute that
$\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(a_{-1})-\tau(\psi))}{\nu_{0}(a_{1})}\cdot\max\{a_{e},A_{e}\}\cdot\dim V_{\beta_{0}}$
$\leq$ $\dim\mathbf{G}-\frac{\tau(\psi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1}).$
Now to derive Theorem 1.6 from Theorems 1.4 and 1.5, it suffices to prove that for any $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$ and $\tau(\psi)\leq\beta_{0}(wa_{-1}w^{-1})$
$\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau(\psi))}{\nu_{0}(a_{1})}\cdot\max\{a_{w},A_{w}\}\cdot\dim V_{\beta_{0}}$
$\leq$ $\dim\mathbf{G}-\frac{\tau(\psi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1}).$
Note that
$\beta_{0}=$ $\kappa\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha$
$=$ $\kappa\cdot\left(\sum_{\alpha\in\Phi(w\mathbf{F}^{w}w^{-1})}\alpha+\sum_{\alpha\in\Phi(w\mathbf{F}_{w}w^{-1})}\alpha\right)$
and
$\beta_{0}(wa_{-1}w^{-1})=\kappa\cdot\left(\sum_{\alpha\in\Phi(\mathbf{F}^{w})}\alpha(a_{-1})+\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\alpha(a_{-1})\right).$
So by Lemma 7.1 and Lemma 7.2, we have
$-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau(\psi))}{\nu_{0}(a_{1})}\cdot\max\{a_{w},A_{w}\}\cdot\dim V_{\beta_{0}}$
$\leq$ $-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{(\beta_{0}(wa_{-1}w^{-1})-\tau(\psi))}{\nu_{0}(a_{1})}\cdot a_{0}\cdot\dim V_{\beta_{0}}$
$=$ $-\sum_{\alpha\in\Phi(\mathbf{F}_{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}+\frac{\beta_{0}(wa_{-1}w^{-1})}{\nu_{0}(a_{1})|\kappa|}-\frac{\tau(\psi)}{\nu_{0}(a_{1})|\kappa|}$
$=$ $\sum_{\alpha\in\Phi(\mathbf{F}^{w})}\frac{\alpha(a_{1})}{\nu_{0}(a_{1})}-\frac{\tau(\psi)}{\nu_{0}(a_{1})|\kappa|}$
$\leq$ $-\frac{\tau(\psi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1}).$
Here we use the fact that all $\mathbb{Q}$-roots in $\Phi(\mathbf{F}^{w})$ are positive and $\alpha(a_{1})\leq 0$ for any $\alpha\in\Phi(\mathbf{F}^{w})$. This completes the proof of the theorem. ∎
Proofs of Theorems 1.7 and 1.8.
@@
where $a=\operatorname{diag}(e^{t_{1}},e^{t_{2}},\cdots,e^{t_{n}})$.
Let $\rho$ be the standard representation of $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$. Let $\{e_{1},e_{2},...,e_{n}\}$ be the standard basis in $V$ where
@@
One may verify that $\beta_{0}=(-\frac{1}{n})\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha$. Theorem 1.7 follows from Theorem 1.6.
Now consider the adjoint representation of $\mathbf{G}=\operatorname{\mathbf{SL}}_{n}$. Let $V_{\beta_{0}}$ be the highest weight space in $\mathfrak{sl}_{n}$
@@
Note that here $\nu_{0}=-\beta_{0}$ is the highest root in $\mathfrak{sl}_{n}$ and $\beta_{0}(a_{t})=(b_{n}-b_{1})t<0$ $(t>0)$. The stabilizer $\mathbf{P}_{\beta_{0}}$ of $V_{\beta_{0}}$ and $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ are